OpenStudy (cindy):

simplify its hard dont understand it at all 2*3+4/2(3+4) need help om this one

OpenStudy (anonymous):

6+4/2(7) 6+(2*7) 6+14 = 20

OpenStudy (cindy):

but wat happens t the 4 in the first step you showed me

OpenStudy (anonymous):

it is divided by 2

OpenStudy (anonymous):

2*3 {6} + [(4/2)(3+4)] {2*7} [6] + [2*7] = 6 + 14 = 20